# Unsupervised Episode Detection for Large-Scale News Events

### v1

@@ 4. Methodology
In order to tackle the episode detection task, we propose a novel unsupervised framework, EpiMine. As shown in Figure 2, EpiMine consists of the following four core components: (1) episode indicative term mining, which identifies combinations of salient terms that are likely to discriminatively co-occur within an episode and not across episodes; (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
4.1. Episode Indicative Term Mining
@@
Figure 3. Natural partition between two episodes within a key event article. An episode’s discriminative terms are bolded, while salient but non-discriminative terms are underlined.
@@
In the absence of supervision, our goal is to identify potential candidates for episodes. Episodes are often described in relation to each other and usually lack specific timestamps or locations consistently mentioned within their segments. For example, the phrase “police dispersed protesters” may not have a precise timestamp because it is a response to “protesters stormed the Legislative Council Complex,” and some journalists may consider the implicit ordering adequate.
Additionally, the same episode can be described using different entities and actions, as journalists may highlight different perspectives when reporting an episode. For example, both “protesters shoved against the barricades” and “the police used pepper-spray on the protesters” describe the episode “protesters stormed the Legislative Council Complex”. However, they are semantically different, focusing on different core entities and actions. Therefore, we cannot depend on a consistent subject-action-object triple or an explicit time and location associated with each episode in the articles.
Salient Terms.
@@
To circumvent this challenge, we exploit the idea that journalists naturally partition news articles according to episodes, forming episode fragments. For example, as shown in Figure 3, an article will likely complete its discussion of episode #1, “Protesters stormed the Legislative Council Complex” (red), before fully shifting to discussing episode #2, “protesters vandalized the Legislative Chamber” (blue). Across these episode fragments, certain salient terms are featured (e.g., protesters, legislative, vandalizing, podium). We modify the idea of event salience from (Jiao et al., 2023) for the task of episode detection:
Definition 4.1 (Salience).
@@
(1) $\textbf{Salience}(w_{i})=\left(1+\log^{2}\left(freq(w_{i})\right)\right)\log\left(\frac{N_{bg}}{bgf(w_{i})}\right)$
Discriminative Co-occurrence.
Now, each text segment within the key event corpus has a corresponding set of salient terms identified. In Figure 3, we can see that the first episode fragment, and Episode #1 in general, features a combination of similar terms, such as “protesters”, “barrier”, and “breach”. Likewise, the second episode may include a combination of terms similar to “protesters”, “spray-painting”, and/or “flag”. Hence, we introduce the significance of co-occurrence for episode detection. Despite some journalists choosing to only describe the protesters spray-painting while others focus on the protesters draping the colonial-era flag, we must be able to recognize that their respective salient terms are likely to co-occur within the same episode.
However, we make a novel distinction between a co-occurrence and a discriminative co-occurrence. A salient term $a$ (e.g., “protesters”) may often co-occur with another salient term $b$ (“spray-painting”) within an episode. On the other hand, if $a$ frequently co-occurs with many other terms in various episodes (“protesters broke”), $a$ and its co-occurrences are less useful for distinguishing episodes. Thus, $(a,b)$ is not a discriminative co-occurrence.
Definition 4.2 (Discriminative Co-occurrence).
@@
We compute the discriminative occurrence $d$ between salient term pair $(a,b)$ using the following equation:
(2) $\begin{split}\textbf{d}(a,b)=&\log\left(\frac{freq(a,b)}{\max(\bar{f_{a}},\bar{f_{b}})}\right)\times\log\left(\frac{|T|}{max(|F_{a}|,|F_{b}|)}\right),\\ \text{where }&\bar{f_{a}}=\frac{1}{|T|}\sum\limits_{\forall w_{i}\in T}freq(a,w_{i})\text{, and}\\ &F_{a}=\{freq(a,w_{i})>1\text{ }\forall\text{ }w_{i}\in T\}\end{split}$
The first $\log$ term ensures that the pair’s co-occurrence ($freq(a,b)$) is statistically significant ($\geq$ the max of $a$ and $b$’s mean vocabulary-wide co-occurrence respectively). The second $\log$ term ensures that the salient term pair is a discriminative matching, penalizing pairings where either $a$ or $b$ co-occurs with a large proportion of the salient term set $T$). For example, co-occurrences with “protesters” are not discriminative because “protesters” is a core entity in all episodes and thus frequently co-occurs with many terms in $T$. In contrast, (“slogans”, “flags”) is a discriminative co-occurrence since both terms frequently appear together in segments discussing episode #2 and rarely co-occur with other terms $w_{i}\in T$. If $a$ and $b$ are the same term or close synonyms (determined by statistically significant semantic similarity), they have maximum co-occurrence. By leveraging multiple articles in a large key event corpus, we have sufficient statistical support to ensure our output reflects the average realistic reporting of the key event and its episodes.
4.2. Episode Partitioning
@@
  1. (1)
    Protesters defaced the emblem of Hong Kong, spray-painted slogans, and unfurled the colonial-era flag.
  2. (2)
    The portrait of LegCo president Andrew Leung was defaced.
  3. (3)
@@
We can naturally see that segments 1-3 all discuss the “protesters vandalized the Legislative Chamber” episode, while segment 4 discusses the “police dispersed protesters” episode. We can systematically replicate this partitioning process by considering the discriminative co-occurrence score between all pairwise combinations of terms from segments ($i-1$) and ($i$). If the average discriminative co-occurrence and static semantic similarity between each term $a$ from ($i-1$) and $b$ from ($i$) is statistically significant ($\geq\mu_{d}-\sigma_{d}$) for that specific article $d$ (e.g., notably (slogans, defaced) for segments 1-3), we hypothesize that the same episode is being discussed and merge them into one episode fragment. If not (e.g., (slogans, arrested) for segments 3-4), this indicates that a different episode is being discussed, and we partition them into two episode fragments. Further implementation details are provided in Appendix B.
4.3. Candidate Episode Estimation
After partitioning each article into episode fragments, we can identify the set of articles that maximizes the quantity and quality of potential episodes. This step is crucial as it enables EpiMine to minimize noise from articles that may discuss only a subset or none of the ground-truth episodes (e.g., focusing on high-level analyses or other relevant key events). Simultaneously, it improves our method’s efficiency by not relying on the entire key event corpus. We compute the per-article rank by multiplying two metrics:
  1. (1)
@@ 4.3. Candidate Episode Estimation
    Quality of episode fragments: A top article should primarily consist of episode fragments containing salient terms that discriminatively co-occur. This reduces the rank of general fragments which summarize/analyze the event. We use the average of each episode fragment’s mean inner-discriminative co-occurrence (across all pairwise combinations of a fragment’s salient terms).
  2. (2)
@@
    Quantity of episode fragments: A top article should ideally contain the highest number of ground-truth episodes. Therefore, we take the $\log$ of the number of episode fragments in the article.
@@
After ranking all articles, we select the top $\delta$to serve as the basis for candidate episode estimation. To estimate the candidate episodes, we must first resolve potential co-references to the same episode across these top articles. We apply agglomerative clustering (Murtagh and Contreras, 2012) to the top episode fragments using a pre-computed distance matrix. The distance between two fragments (inversed) is calculated using the same discriminative and static semantic similarity score used in Section 4.2). We then prune clusters that do not contain a statistically significant number of episode fragments.
@@
Finally, we use an LLM to resolve (1) the lack of time and location stamps in an episode fragment and (2) semantic inconsistency between the fragments within a cluster. The LLM leverages its reasoning capabilities to estimate the candidate episodes by identifying the core attributes of each unique cluster (entities, actions, objects, location, and time). It outputs the episode attributes, relevant keywords (assists with extraction), and top extracted text segments based on the clusters (prompt & example in Appendix C).
4.4. Episode-Segment Classification
With these core summaries of the episode clusters, we obtain a generalized description of each candidate episode. This allows us to compute a simple episode representation for assigning an episode and confidence score to each input segment.
Episode & Segment Representations.
@@ 4.4. Episode-Segment Classification
We compute a candidate episode representation by encoding both its LLM core attributes and extracted segments using SentenceTransformers (ST) (Reimers and Gurevych, 2019). Following extremely weakly supervised text classification works, such as (Wang et al., 2021; Kargupta et al., 2023), we take the harmonic mean of these representations, as the latter extracted segments are likely not as significant as the earlier extractions and core attributes. We similarly encode all input segments with the same ST model.
Episode-Segment Confidence Estimation
Following recent work on extremely weakly supervised text classification (Kargupta et al., 2023), we compute a confidence score specific to the episode-segment classification task. Simply mapping a text segment to its top episode (based on highest cosine similarity) overlooks the presence of episode-irrelevant text segments and fails to address segments discussing multiple episodes (e.g., a journalist’s key event summary at the beginning of an article). We aim to avoid assigning such segments to any cluster, ensuring no overlap between episode clusters, as discussed in Section 3.1.
@@
Thus, we compute segment $s_{i}$’s cosine similarity to its top two episodes ($e_{i}^{0}$ and $e_{i}^{1}$). A larger gap ($e_{i}^{0}-e_{i}^{1}$) between the two indicates higher confidence in mapping $s_{i}$ to its top episode. We normalize each gap by the sum of all segment episode gaps within the entire corpus, ensuring the confidence is relative to the key event:
(3) $s_{i,\text{confidence}}=\frac{e_{i}^{0}-e_{i}^{1}}{\sum_{l=1}^{|\mathcal{S}|}(e^{0}_{l}-e^{1}_{l})}$
@@
After computing each segment’s confidence in its top episode, we assign only those segments with a statistically significant confidence level to their respective episode clusters $E$. We then prune candidate episode clusters with $\leq 0$ assigned segments, resulting in our final detected episodes and their clusters, $\mathcal{E}$.
5. Experimental Setup
For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$. All other hyperparameters are set to their respective default values. To determine statistical significance, we check for whether or not a value is $\geq\mu-\sigma$. For our word representations, we use bert-base-uncased. For our sentence representations, we use all-mpnet-base-v2. We choose to use Claude-2.1222claude.ai/ for fluent candidate episode estimation due to its strong structured JSON/XML input and output formatting abilities. However, this proprietary model can be replaced with any open-source model as EpiMine is model-agnostic. We use only one NVIDIA GeForce GTX 1080 for all experiments.
Table 1. Statistics of our collected datasets. The numbers are averaged per key event.
@@ 5. Experimental Setup
Table 2. Results averaged across each theme, including the mean # of episodes EpiMine identifies per theme. Results are computed on each key event corpus using the top-5 documents for each detected episode. Due to variance in LLM generation, we run it 10 times and report the average of each measure. Bold values denote the top method; $\dagger$ denotes the second-best method.
Terrorism (avg. # of eps = 5.36) Natural Disasters (avg. # of eps = 7.4) Politics (avg. # of eps = 7.5)
@@
EMiner 0.0864 0.0025 0.0048 0.1037 0.0019 0.0037 0.0866 0.0016 0.0032
K-means 0.2123 0.2123 0.2123 0.2785 0.2847 0.2814 0.1604 0.1604 0.1604
K-means + LLM 0.2858 0.1404 0.1826 0.3740 0.1658 0.22 0.2718 0.1736 0.1825
EvMine 0.2303 0.1502 0.1745 0.2815 0.0802 0.1225 0.0536 0.04 0.0458
EvMine + LLM 0.3788 0.1570 0.2133 0.4356 0.1322 0.1940 0.3273 0.1298 0.1728
EpiMine 0.7121 0.2207 0.3243 0.7098 0.2846 0.3453 0.6267 0.2154 0.2923
- No Confidence 0.6197 0.3019 0.3845 0.4366 0.2078 0.2776 0.6029 0.2773 0.3385
Table 3. Ablation studies conducted on top 25% of article episode clusters (Section 4.3).

### v2

@@ 4 Methodology
To tackle episode detection, we propose a novel unsupervised framework, EpiMine. As shown in Figure 2, EpiMine consists of the following four core components: (1) episode indicative term mining, which identifies combinations of salient terms likely to discriminatively co-occur within an episode and not across episodes; (2) episode partitioning, which partitions each article into approximate isolated episodes based on consecutive shifts in the term co-occurrence distribution, (3) LLM-enhanced candidate episode estimation, which clusters the top partitions into candidate episodes and utilizes LLM-based reasoning to produce fluent and meaningful episodes, and (4) episode-segment classification, which maps confident segments to their respective episode clusters.
4.1 Episode Indicative Term Mining
@@
Figure 3: Natural partition between two episodes in a key event article. An episode’s discriminative terms are bolded; salient non-discriminative terms are underlined.
@@
Lacking supervision, our goal is to identify potential candidates for episodes. Episodes are often described in relation to each other and usually lack timestamps or locations consistently mentioned within their segments. For example, the phrase “police dispersed protesters” may not have a precise timestamp because it is a response to “protesters stormed the Legislative Council Complex,” and some journalists may consider the implicit ordering adequate. Additionally, the same episode can be described using different entities and actions— journalists may report different perspectives. For example, both “protesters shoved against the barricades” and “the police used pepper-spray on the protesters” describe the episode “protesters stormed the Legislative Council Complex”. However, they are semantically different, focused on different core entities and actions. Thus, we cannot depend on a consistent subject-action-object triple or an explicit time/location mapped to each episode in the article.
@@
To circumvent this challenge, we exploit the idea that journalists naturally partition news articles according to episodes, forming episode fragments. For example, as shown in Figure 3, an article will likely complete its discussion of episode #1, “Protesters stormed the Legislative Council Complex” (red), before fully shifting to discussing episode #2, “protesters vandalized the Legislative Chamber” (blue). Across these episode fragments, certain salient terms are featured (e.g., protesters, legislative, vandalizing, podium). We adapt the idea of event salience from Jiao et al. (2023) specifically for the task of episode detection.
Definition 3 (Salience).
@@
Thus, we identify a set of salient terms for each segment within the corpus (salience score details in Appendix E).
Discriminative Co-occurrence. In Figure 3, we can see that the first episode fragment, and Episode #1 in general, features a combination of similar terms, such as “protesters”, “barrier”, and “breach”. Likewise, the second episode may include a combination of terms similar to “protesters”, “spray-painting”, and/or “flag”. We note that despite some journalists choosing to only describe the protesters spray-painting, while others focus on the protesters draping the colonial-era flag, we must be able to recognize that their respective salient terms are likely to co-occur within the same episode.
However, we make a novel distinction between a co-occurrence and a discriminative co-occurrence. Salient terms $a$ and $b$ (e.g., “protesters” and “spray-painting”) may often co-occur within an episode. However, if $a$ also frequently co-occurs with many other terms in various episodes (“protesters broke”), $a$ and its co-occurrences are less useful for distinguishing episodes. Thus, $(a,b)$ is not a discriminative co-occurrence.
Definition 4 (Discriminative Co-occurrence).
@@
We compute the discriminative occurrence $d$ between salient term pair $(a,b)$ using the following:
$\begin{split}\textbf{d}(a,b)=&\log\left(\frac{freq(a,b)}{\max(\bar{f_{a}},\bar{f_{b}})}\right)\times\log\left(\frac{|T|}{max(|F_{a}|,|F_{b}|)}\right),\\ \text{where }&\bar{f_{a}}=\frac{1}{|T|}\sum\limits_{\forall w_{i}\in T}freq(a,w_{i})\text{, and}\\ &F_{a}=\{freq(a,w_{i})>1\text{ }\forall\text{ }w_{i}\in T\}\end{split}$ (1)
The first $\log$ term ensures that the pair’s co-occurrence ($freq(a,b)$) is statistically significant ($\geq$ the max of $a$ and $b$’s mean vocabulary-wide co-occurrence respectively). The second $\log$ term ensures the pair is a discriminative match, penalizing cases where $a$ or $b$ frequently co-occurs with a large portion of the salient term set $T$. For example, co-occurrences with “protesters” are not discriminative because “protesters” is a core entity in all episodes and thus frequently co-occurs with many terms in $T$. In contrast, (“slogans”, “flags”) is a discriminative co-occurrence since both terms frequently appear together in segments discussing episode #2 and rarely co-occur with other terms $w_{i}\in T$. If $a$ and $b$ are the same term or close synonyms (determined by statistically significant semantic similarity), they have maximum co-occurrence. By leveraging multiple articles in a large key event corpus, we have sufficient statistical support to ensure our output reflects the average realistic reporting of the key event and its episodes.
4.2 Episode Partitioning
@@
  1. 1.
    Protesters defaced the Hong Kong emblem, spray-painted slogans, and unfurled the flag.
  2. 2.
    The portrait of LegCo president was defaced.
  3. 3.
@@
We can naturally see that segments 1-3 all discuss the “protesters vandalized the Legislative Chamber” episode, while segment 4 discusses the “police dispersed protesters” episode. We can systematically replicate this partitioning process by considering the discriminative co-occurrence score between all pairwise combinations of terms from segments ($i-1$) and ($i$). If the average discriminative co-occurrence and static semantic similarity between each term $a$ from ($i-1$) and $b$ from ($i$) is statistically significant ($\geq\mu_{d}-\sigma_{d}$) for that specific article $d$ (e.g., notably (slogans, defaced) for segments 1-3), we hypothesize that the same episode is being discussed and merge them into one episode fragment. If not (e.g., (slogans, arrested) for segments 3-4), this indicates that a different episode is being discussed, and we partition them into two episode fragments. Further implementation details are provided in Appendix G.
4.3 LLM-Enhanced Episode Estimation
LLMs demonstrate strong event-specific reasoning at the phrase or sentence level (Pai et al., 2024; Gao et al., 2024), but they struggle with understanding long contexts (Li et al., 2024; Liu et al., 2024). This limitation hampers their ability to process all episode fragments for detecting episodes. Additionally, noisy retrieval significantly affects reasoning performance (Shen et al., 2024). To address these challenges, we propose a synergistic approach that enhances in-context episode reasoning by reducing the number of required fragments while improving their cohesiveness and quality. We first identify the set of articles that maximizes the quantity and quality of potential episodes, where each article is ranked by multiplying two metrics:
  1. 1.
@@ 4.3 LLM-Enhanced Episode Estimation
    Quality of episode fragments: A top article should primarily consist of episode fragments containing salient terms that discriminatively co-occur. This reduces the rank of general fragments which summarize/analyze the event. We average each episode fragment’s mean inner-discriminative co-occurrence (across all pairwise combinations of its salient terms).
  2. 2.
@@
    Quantity of episode fragments: A top article should ideally contain all ground-truth episodes. Therefore, we take the $\log$ of the number of episode fragments in the article.
@@
After ranking all articles, we select the top $\delta$and resolve potential co-references to the same episode across these top articles. We apply agglomerative clustering Murtagh and Contreras (2012) to the top episode fragments using a pre-computed distance matrix. The distance between two fragments (inversed) is calculated using the same discriminative and static semantic similarity score used in Section 4.2). Clusters with a statistically insignificant number of episode fragments are pruned.
@@
Finally, we provide episode fragment clusters as a more interpretable context for the LLM to resolve two challenges: (1) missing time and location stamps in fragments, and (2) semantic inconsistencies within clusters. The LLM summarizes each cluster by identifying its core attributes— entities, actions, objects, location, and time. It then outputs the episode attributes, relevant keywords for extraction, and the top extracted text segments (prompt & example in Appendix H).
4.4 Episode-Segment Classification
@@
With these core summaries of the episode clusters, we obtain a generalized description of each candidate episode. For each candidate, we encode its LLM-based core attributes and extracted segments to compute a simple episode representation. Specifically, following extremely weakly supervised text classification works Wang et al. (2021); Kargupta et al. (2023), we take the harmonic mean of these representations—as the latter extracted segments are likely not as significant as the earlier extractions and core attributes. We similarly encode all input segments with the same encoder. We use these to assign an episode and confidence score to each encoded input segment.
Episode-Segment Confidence Estimation.
Directly mapping a text segment to its top episode based on cosine similarity risks misclassifying episode-irrelevant segments or those discussing multiple episodes (e.g., a journalist’s summary). To avoid classifying such segments and ensure non-overlapping episode clusters (as discussed in Section 3.1), we must determine the confidence of a segment discussing a single episode.
@@
We compute segment $s_{i}$’s cosine similarity to its top two episodes ($e_{i}^{0}$ and $e_{i}^{1}$). A larger gap ($e_{i}^{0}-e_{i}^{1}$) reflects greater confidence in classifying $s_{i}$ to $e_{i}^{0}$. Each gap is normalized by the sum of all segment-episode gaps across the corpus, ensuring confidence is relative to the key event:
$s_{i,\text{confidence}}=\frac{e_{i}^{0}-e_{i}^{1}}{\sum_{l=1}^{|\mathcal{S}|}(e^{0}_{l}-e^{1}_{l})}$ (2)
@@
Segments with statistically significant confidence in their top episode are assigned to their respective episode clusters $E_{i}$. Episodes with no assigned segments are pruned, yielding the final detected episodes and clusters, $\mathcal{E}$.
5 Experiments
For implementing EpiMine, we use the following hyperparameters across all datasets: $\delta=25\%$, $sim\_thresh=0.75$. We also use Claude-2.1 as our base LLM ([Uncaptioned image]). All other hyperparameters are set to their respective default values. We provide all experimental settings in Appendix A.
Table 1: Statistics of our collected datasets. The numbers are averaged per key event.
@@ 5 Experiments
Table 2: Results averaged across each theme, including the mean # of episodes that EpiMine identifies per theme (in parenthesis). Results are computed on each key event corpus using the top-5 documents for each detected episode. Due to variance in LLM generation, we run it 10 times and report the average of each measure. We scale each value by 100. Bold values denote the top method; second-best method is underlined.
Terrorism (5.36 eps) Natural Disasters (7.4 eps) Politics (7.5 eps)
@@
EMiner 8.64 0.25 0.48 10.37 0.19 0.37 8.66 0.16 0.32
K-means 21.23 21.23 21.23 27.85 28.47 28.14 16.04 16.04 16.04
K-means + [Uncaptioned image] 28.58 14.04 18.26 37.40 16.58 22.00 27.18 17.36 18.25
EvMine 23.03 15.02 17.45 28.15 8.02 12.25 5.36 4.00 4.58
EvMine + [Uncaptioned image] 37.88 15.70 21.33 43.56 13.22 19.40 32.73 12.98 17.28
EpiMine ([Uncaptioned image]) 71.21 22.07 32.43 70.98 28.46 34.53 62.67 21.54 29.23
- No Confidence ([Uncaptioned image]) 61.97 30.19 38.45 43.66 20.78 27.76 60.29 27.73 24.77
- No LLM 37.73 21.62 24.77 37.19 14.78 17.52 30.64 23.51 19.06
5.1 Datasets
We conduct our experiments on three novel thematic, real-world news corpora selected from Wikipedia [^2] https://en.wikipedia.org/wiki/ over the last decade. For each theme, we manually collect approximately $10$ key events composed of multiple articles and ensure that distinct episodes exist in each:
  • Natural disasters: 2023 Tornado outbreak sequence; 2023 Hawaii Wildfires; 2021 Western Kentucky tornado; 2017 Mocoa landslide; 2010 Haiti earthquake; 2021 Henan floods; 2019 Nyonoksa radiation accident; 2022 NA winter storm; 2011 Fukushima nuclear accident
  • Political events: 2020 Kyrgyz Revolution, 2019 Storming of the Hong Kong Legislative Council Complex; 2019 Siege of the Hong Kong Polytechnic University; 2017 Zimbabwean coup; 2018 Italian government formation; 2021 January 6 U.S. Capitol attack; 2018 Thai Cave Rescue Operation; 2018 Armenian Revolution; 2017 Lebanon–Saudi Arabia dispute; 2013 Tunisian political crisis
The articles are obtained from the Wikipage references of each key event— filtered with constraints in time, language, and relevance. Furthermore, each article is segmented to match our setting (Appendix F), and each segment is automatically annotated (Appendix I) with either its corresponding episode ID or with a multiple/no episode tag (‘M’ or ‘X’). Further details on the criteria/process, each theme, and corresponding key events are in Appendices D and I. We also conduct a human-automatic agreement analysis for segment-episode annotation, which shows substantial agreement with good reliability (Appendix D).
5.2 Baselines
We compare against the following methods using the evaluation metrics specified in Appendix C: (1) K-means Likas et al. (2003): given the # of ground-truth episodes, it clusters segments using ST  Reimers and Gurevych (2019) embeddings; (2) EvMine Zhang et al. (2022): a document-level unsupervised key event detection method adapted to segment level for episode detection; (3) EMiner Jiao et al. (2023): unsupervised event chain miner that clusters atomic actions, adapted to episodes; (4) No Confidence: an ablation that uses max cosine-similarity instead of confidence from Equation 2; (5) No LLM: an ablation that uses estimated episode clusters from Section 4.3 to compute our episode representations directly. We also integrate [Uncaptioned image] into K-means and EvMine using our same prompt (Appendix H). All baseline and ablation details are in Appendix B.
Table 3: Ablation studies conducted on top 25% of article episode clusters (Section 4.3).